Let f(x)=1-2x and find
a) f^-1 (x) and b) 1/f(x)

Answer by edjones(8007)

f(x)=1-2x
a)
y=1-2x
x=1-2y switch x and y, then solve for y.
-2y=x-1
y=(x-1)/-2
f^-1(x)=(1-x)/2
.
b)
1/f(x)=1/(1-2x)
.
This proves that an inverse is not the same as a reciprocal.
.
Ed